# Key Words.: Galaxies: clusters: general – Galaxies: clusters: individual: Abell 426 – X-rays: galaxies: clusters – Radio continuum: galaxies – Galaxies: clusters: intracluster medium – Galaxies: kinematics and dynamics1 Introduction

### v1

@@ Appendix A Chandra and XMM datasets
For Chandra:
11713, 11714, 11715, 11716, 12025, 12033, 12036, 12037, 13989, 13990, 13991, 13992, 3209, 4289, 4946, 4947, 4948, 4949, 4950, 4951, 4952, 4953, 6139, 6145, 6146.
The initial data processing was done following the procedures described in Vikhlinin et al. (2005).
For XMM-Newton: 0085110101, 0085590201, 0151560101, 0204720101, 0204720201, 0305690101, 0305690301, 0305690401, 0305720101, 0305720301, 0305780101, 0405410101, 0405410201, 0673020201, 0673020301, 0673020401. The XMM-Newton data were cleaned for flares and then used similarly to Churazov et al. (2003).
Appendix B Radial X-ray surface brightness profile
@@
As described in Sect. 2.3, eROSITA, Chandra, and XMM-Newton images were rebinned to the same spatial grid (2” pixels) and normalized to ensure a common flux calibration. A radial profile of the resulting image is shown in Fig. 12. The observed profile was approximated with a simple combination of a -model (the blue line in Fig. 12) and a constant (the dashed line). The very core (”) was excluded when fitting the profile. This model (sum of the -model and the constant) is used to remove a symmetric component from the X-ray image and express the residual deviations relative to the same model.

### v2

@@ Appendix A Chandra, XMM-Newton, and SRG/eROSITA datasets
For Chandra: 11713, 11714, 11715, 11716, 12025, 12033, 12036, 12037, 13989, 13990, 13991, 13992, 3209, 4289, 4946, 4947, 4948, 4949, 4950, 4951, 4952, 4953, 6139, 6145, 6146.
Observations were processed with the standard Chandra data reduction (CIAO v. 4.17) and calibration software (CalDB v. 4.12). Data analysis steps are described in detail in Vikhlinin et al. (2009) and include high background period filtering, application of the latest calibration corrections to the detected X-ray photons, and determination of the background intensity in each observation. In practice, for the imaging analysis of the central region of the Perseus cluster, the details of the background calibration are not important.
For XMM-Newton: 0085110101, 0085590201, 0151560101, 0204720101, 0204720201, 0305690101, 0305690301, 0305690401, 0305720101, 0305720301, 0305780101, 0405410101, 0405410201, 0673020201, 0673020301, 0673020401.
The XMM-Newton data were cleaned for flares and then processed similarly to Churazov et al. (2003), utilizing recent versions of the calibration files. Only the MOS data were used. The steady component of the detector background was subtracted using data from blank field observations. The variable component of the detector background was modeled as a power law as a function of energy . The data in the 11-12 keV range were used to calculate the normalization of this variable component. As with the Chandra data, only the central part of the cluster measured by XMM contributes to the combined image, and, therefore, the results are not affected by the treatment of the background.
For eROSITA: data from all four consecutive scans are combined after filtering for the periods of enhanced solar activity, which cause short periods of strongly elevated levels of the instrumental background. For the imaging analysis, the data taken with all seven telescope modules (TMs) are combined. Data reduction, filtering, vignetting-correction, and background subtraction are performed in the same way as was done in the previous studies exploring Galactic diffuse X-ray sources, including those with the surface brightness below the total (sky and detector) background level (Churazov et al. 2021b; Khabibullin et al. 2022, 2023, 2024). Specifically, the energy-dependent contribution of the instrumental background is modeled and subtracted based on the calibration data accumulated via observations with the ”closed filter wheel” configuration, while corrections for exposure time and vignetting are conducted so that the data are characterized by field-of-view-averaged response matrices. Deep observations of the Coma cluster (Churazov et al. 2021a) were instrumental for eROSITA’s spectral cross-calibration via ICM temperature measurement in its core, while sets of bright point sources observed during the all-sky survey were used to construct the shape of the off-axis PSF (Churazov et al. 2023).
Appendix B Radially dependent weights
As described in Sect. 2.3, eROSITA, Chandra, and XMM-Newton images were rebinned to the same spatial grid (2” pixels) and re-normalized to ensure a common flux calibration. For this re-normalization, a circle ( radius) centered at NGC1275 was used, and all fluxes were reduced to the eROSITA values. The last step was the coadding of images with radially dependent weights
where is the image’s indices (1,2,3 for Chandra, XMM-Newton, and eROSITA, respectively). The weights are defined as
where is the distance from NGC1275 in arcminutes; for eROSITA, Chandra, and XMM, respectively, and .
The main goal of this weighting scheme is to preserve Chandra’s high angular resolution in the core and suppress the noisy contributions at the edges of Chandra and XMM-Newton images. The implemented weights are shown in Fig. 12. The individual images are shown in Fig. 13.
Appendix C Radial X-ray surface brightness profile
@@
A radial profile of the resulting image is shown in Fig. 14. The observed profile was approximated with a simple combination of a -model (the blue line in Fig. 14) and a constant (the dashed line). The very core (”) was excluded when fitting the profile. This model (sum of the -model and the constant) is used to remove a symmetric component from the X-ray image and express the residual deviations relative to the same model.
Appendix D Impact of photoelectric absorption in the Milky Way on the surface brightness variations.
The center of the Perseus cluster is only degrees away from the Galactic Plane. This means that some of the observed variations of the X-ray surface brightness might be caused by spatially non-uniform photoelectric absorption of X-rays by neutral or molecular gas in the Galaxy. Here, we demonstrate that the amplitude of this effect is significantly smaller than the observed variations. To do so, we used the 3D dust maps from Bayestar19 (see Green et al. 2019, for detailed description). We are looking for the intervening absorption of extragalactic X-rays from the cluster itself and from the Cosmic X-ray Background (CXB) in general. The total column density (NH) was calculated from the extinction maps based on the Bayestar19 data. The column density map was converted into an ”attenuation map” of the X-ray flux in the 0.4-2.3 keV band, taking into account the eROSITA response. As a final step, the attenuation map in every pixel was divided by the mean value in concentric rings, mimicking the procedure used to generate Fig. 4. A direct comparison of the expected modulation with the X-ray data (Fig. 15) shows that the photoelectric absorption effect is subdominant. Furthermore, we have verified that in the harder 1-2.3 keV band, where photoelectric absorption is weaker, the X-ray image possesses similar large-scale structures as seen in Fig. 6.